### v3

# thanks: corresponding authorthanks: corresponding author

Abstract
Post-selection strategies that discard low-confidence computational results can significantly improve the effective fidelity of quantum error correction at the cost of reduced acceptance rates, which can be particularly useful for offline resource state generation. Prior work has primarily relied on the “logical gap” metric with the minimum-weight perfect matching decoder, but this approach faces fundamental limitations including computational overhead that scales exponentially with the number of logical qubits and poor generalizability to arbitrary codes beyond surface codes. We develop post-selection strategies based on computationally efficient heuristic confidence metrics that leverage error cluster statistics (specifically, aggregated cluster sizes and log-likelihood ratios) from clustering-based decoders, which are applicable to arbitrary quantum low-density parity check (QLDPC) codes. We validate our method through extensive numerical simulations on surface codes, bivariate bicycle codes, and hypergraph product codes, demonstrating orders of magnitude reductions in logical error rates with moderate abort rates. For instance, applying our strategy to the $[[144,12,12]]$ bivariate bicycle code achieves approximately three orders of magnitude reduction in the logical error rate with an abort rate of only 1% (19%) at a physical error rate of 0.1% (0.3%). Additionally, we integrate our approach with the sliding-window framework for real-time decoding, featuring early mid-circuit abort decisions that eliminate unnecessary overheads. Notably, its performance matches or even surpasses the original strategy for global decoding, while exhibiting favorable scaling in the number of rounds. Our approach provides a practical foundation for efficient post-selection in fault-tolerant quantum computing with QLDPC codes.
1 Introduction
Quantum computing promises exponential speedups for certain computational problems, but realizing this potential requires overcoming the fundamental challenge of quantum decoherence and operational errors [1]. Quantum error correction (QEC) provides a path toward fault-tolerant quantum computing by encoding logical qubits into larger systems of physical qubits, enabling the detection and correction of errors during computation. However, QEC typically requires substantial resources, posing a major challenge for its realization on physical hardware.
A promising technique to mitigate the resource overhead of QEC is post-selection, which strategically discards low-confidence computations, thereby achieving significantly higher reliability from the remaining accepted results. This approach enables attaining very low logical error rates with relatively small codes, with the trade-off of non-determinism.
Post-selection can be particularly useful for offline resource state generation processes such as magic state preparation [2, 3], where aborting and retrying failed attempts does not damage existing encoded quantum information. Additionally, when estimating expectation values of observables (e.g., ground-state energies via the variational quantum eigensolver [4, 5]), discarding low-confidence runs can substantially improve estimator fidelity at the cost of additional sampling in the context of quantum error mitigation [6, 7, 8, 9, 10, 11].
In terms of QEC, initial approaches to post-selection were simple: abort the computation whenever any error was detected through syndrome measurements [12, 13]. This error-detection-based strategy could substantially improve effective error thresholds, but suffers from significant retry overhead costs.
Motivated by these initial works, QEC researchers have explored more sophisticated “partial” post-selection strategies that abort conditionally based on specific criteria even when errors are detected. Namely, decoders produce soft outputs (such as likelihoods) besides final corrections, which may contain information on the reliability of the decoded outcomes and thus can be employed to establish such criteria. For example, an experimental work [14] demonstrated a reduction in the logical error rate by aborting when a correction from a decoder contains ambiguous faults.
The logical gap (defined as the log-likelihood ratio difference between candidate corrections in distinct logical classes) has emerged as a Bayesian-motivated metric quantifying decoding confidence [15, 16, 17, 18]. Physically, according to the statistical-mechanical mapping of error correction, the logical gap of a surface code in the code-capacity setting is proportional to the zero temperature free energy cost of a domain wall in the mapped random-bond Ising model [19, 20]. Post-selection strategies based on the logical gap proved remarkably effective, enabling about 15-fold reductions in error rates with a relative overhead factor of <2<2 [15], error threshold improvements (up to 50% with full post-selection under code capacity noise) [17], and order-of-magnitude reductions in magic state preparation costs [21, 22]. The logical gap has become the standard approach for confidence-based post-selection in QEC.
Nevertheless, despite these advances, the logical gap method faces fundamental limitations that restrict broader applicability. First, computational overhead scales exponentially with the number $k$ of logical qubits (requiring comparative decoding across all $2^{k}$ logical classes), which makes the method prohibitive for codes or circuits involving multiple logical qubits. Second, the method is not guaranteed to work well for general decoders beyond the optimal degenerate maximum likelihood decoder [23, 24] or the minimum-weight perfect matching (MWPM) decoder [25, 26, 27], as a correction within a fixed logical class may fail to represent the likelihood of the class.
Recent efforts have sought to address some of these limitations through alternative strategies, including decoder-independent confidence metrics based on syndrome density [28] and cluster-geometry-based soft outputs [29]. However, these approaches still face challenges: the syndrome density method, while efficient and generalizable across code families, has sub-optimal performance, and the cluster-geometry-based method, while overcoming the exponential overhead problem of the logical gap, is currently applicable only to the repetition and surface codes.
In this work, we develop computationally efficient heuristic confidence metrics that overcome the limitations of the logical gap method, while maintaining effective post-selection performance across diverse quantum low-density parity check (QLDPC) codes. Specifically, we quantify decoding confidence based on the clustering structure of errors revealed by clustering-based decoders, such as the union-find decoder [30, 31], the belief propagation plus localized statistics decoder (BP+LSD) [32], and the ambiguity clustering decoder [33]. Leveraging the intuition that the size and distribution of error clusters directly correlate with difficulty in decoding, we introduce two families of heuristic confidence metrics: the cluster size norm fraction and the cluster log-likelihood ratio (LLR) norm fraction. Unlike the logical gap, our approach requires only a single decoder execution and is applicable to general QLDPC codes.
We demonstrate the effectiveness of our cluster-based confidence metrics through extensive numerical simulations on surface codes [34, 19], bivariate bicycle codes [35, 36], and hypergraph product codes [37, 38, 39] using the BP+LSD decoder. The results clearly show orders of magnitude reduction in logical error rates at moderate abort rates for all the three code families. While our strategy does not match the logical gap method for surface codes (unsurprising given its strong probability-theoretical foundation), our approach’s efficiency and generalizability make it a compelling alternative in practice.
Furthermore, we develop a real-time post-selection strategy based on the sliding-window decoding framework [19, 40, 41, 42, 43, 44, 45, 32, 46], which enables mid-circuit abort decisions that eliminate unnecessary overheads. This strategy exhibits performance that is comparable to (or even better than) the global strategy in terms of the trade-off relation between the logical error rate and the average time cost. Moreover, unlike the global strategy, our real-time strategy maintains a consistent per-round logical error rate and abort rate, which is a favorable property for circuits with large depths.
The main contributions of this work are:
  • We introduce cluster-based confidence metrics that overcome the computational and applicability limitations of the logical gap method, enabling efficient post-selection for general QLDPC codes.
  • We develop a real-time post-selection strategy that integrates naturally with sliding-window decoding, allowing for early abort decisions and reduced computational overhead.
  • We provide comprehensive numerical evidence demonstrating the effectiveness of our approach across multiple code families, achieving substantial reductions in logical error rates with modest abort rates.
2 Results
2.1 Soft outputs from decoding
The minimum requirement of a decoder is to infer whether each logical observable needs to be flipped, given the detector outcomes. Here, detectors indicate specific products of measurement outcomes deterministic in the absence of errors, which are normally constructed from two consecutive measurements of a stabilizer. A detector error model (DEM) [47] specifies the prior probabilities of independent fault locations (referred to as “error mechanisms”) as well as the detectors and logical observables flipped by each of them. Decoders primarily aim to address the classification problem: given a DEM and detector outcomes, which logical class the current state belongs to (i.e., which logical observables have been flipped due to errors). In practice, however, decoders typically produce additional information beyond this basic classification. These additional outputs are referred to as the soft outputs of the decoder.
One common and straightforward soft output is the log-likelihood weight of the correction, defined as the summation of the prior log-likelihood ratios (LLRs) for individual error mechanisms in the correction obtained from the decoder. This approach is possible for many decoders, which return not only the logical class but also an explicit error correction (candidate error configuration) within the class. Specifically, if $p_{e}$ is the prior probability of each error mechanism $e$, the log-likelihood weight of a correction $\tilde{E}$ is given by
$w(\tilde{E})\coloneqq\sum_{e\in\tilde{E}}\log\frac{1-p_{e}}{p_{e}}.$ (1)
In particular, the MWPM decoder [25, 26, 27] identifies a correction that minimizes this log-likelihood weight, which corresponds to a maximum likelihood solution when degeneracy is not considered.
Other useful soft outputs include posterior LLRs of individual error mechanisms for given detector outcomes, which can be estimated using the belief propagation (BP) decoder [48]. Due to inherent degeneracy of QEC codes, BP sometimes fails to converge and therefore does not exhibit an error threshold [49, 50]. Nevertheless, the posterior LLRs can be used in subsequent post-processing routines such as the ordered statistics decoder (OSD) [51, 49] or for running other decoders such as MWPM with updated weight information [52].
Here we focus on a specific application of soft outputs: quantifying decoding confidence. Specifically, we aim to estimate the reliability of the logical flip inferred by a decoder through analysis of its soft outputs. Given such a confidence metric, we construct a post-selection strategy that accepts a trial only when the metric value exceeds a specified cutoff threshold. We note that soft outputs have found broader applications beyond post-selection, including hierarchical code architectures where confidence information from inner codes guides the decoding of outer codes [16, 29, 53]. However, this work focuses specifically on their application to post-selection.
2.1.1 Logical gap for quantifying decoding confidence
A prominent approach in the literature for quantifying decoding confidence involves comparative decoding and evaluation of the logical gap (also known as the complementary gap) [15, 16, 17]. Here, comparative decoding refers to a technique that performs decoding within each logical class and selects the minimum-weight correction among all classes. Decoding within a fixed logical class (identified by a vector $\overrightarrow{\mathbf{\lambda}}\in\mathbb{Z}_{2}^{k}$ for the number $k$ of observables) can be implemented by adding $k$ new “virtual” detectors (corresponding to the $k$ observables) into the DEM and assigning specific values (elements of $\overrightarrow{\mathbf{\lambda}}$) to them for decoding. The logical gap $\Delta$ is then defined as the difference between the two smallest log-likelihood weights among the corrections $\{\tilde{E}_{\overrightarrow{\mathbf{\lambda}}}\}_{\overrightarrow{\mathbf{\lambda}}\in\mathbb{Z}_{2}^{k}}$ in different logical classes, i.e.,
$\Delta\coloneqq w_{(2)}-w_{(1)},$
where $w_{(1)}\leq w_{(2)}\leq\cdots\leq w_{(2^{k})}$ are the sorted values of $w(\tilde{E}_{\overrightarrow{\mathbf{\lambda}}})]_{\overrightarrow{\mathbf{\lambda}}\in\mathbb{Z}_{2}^{k}}$. A large logical gap indicates that error patterns in alternative logical classes are significantly less likely than the predicted one, thereby implying high confidence in the decoding outcome.
Despite this general description, the logical gap method has been investigated almost exclusively for the MWPM decoder. This limitation arises from several fundamental problems that prevent broader applicability. First, unless the MWPM decoder is used, we cannot guarantee that decoding within a fixed logical class yields the minimum-weight correction in that class. This means that the weight may not adequately represent the likelihood of the logical class, which could diminish the effectiveness of the logical gap in capturing decoding confidence. Moreover, some decoders may exhibit significantly degraded performance when the logical class is fixed. For example, the union-find (UF) decoder [30] may not be well-suited for comparative decoding, as a cluster that touches a virtual detector (which connects to $d$ or more errors) can grow abnormally fast, thereby failing to accurately capture the underlying clustering structure of the error pattern. Similarly, decoders for QLDPC codes that rely on the low-density property (such as the generalized UF decoder [31]) may also perform poorly due to high-degree virtual detectors with $O(d)$ connections.
While the MWPM decoder does not suffer from these issues, it is applicable only to a limited class of QEC codes, such as repetition and surface codes, where each elementary error mechanism affects at most two detectors [54]. (DEMs may contain error mechanisms affecting more than two detectors, but they can be decomposed into elementary error mechanisms.) A notable exception is the class of two-dimensional color codes [55], which themselves are not compatible with MWPM but allow suboptimal decoders that operate by applying MWPM to specific matchable subgraphs [56, 57, 58, 59, 60, 61, 62]. This makes the logical gap method reasonably effective [22], particularly with the concatenated MWPM decoder [62].
Another significant limitation of the logical gap method is its computational cost, which scales exponentially with the number $k$ of observables, requiring decoding across all $2^{k}$ logical classes. While this scaling may be manageable for memory experiments involving a single logical qubit, it renders the method essentially impractical for scenarios involving multiple logical qubits, e.g., when using codes that encode multiple logical qubits per block or when multiple logical qubits are subject to correlated noise due to multi-qubit gates.
2.2 Post-selection strategy based on error cluster statistics
Due to the aforementioned limitations of the logical gap method, we seek alternative approaches for quantifying decoding confidence and constructing post-selection strategies. This requires decoders that provide rich soft outputs about error configurations, a condition that clustering-based decoders satisfy well. These decoders operate by inferring the clustering structure of errors (the pattern of how errors group together) which could directly correlate with the difficulty of the decoding process.
2.2.1 Clustering-based decoders
Figure 1: Overview of our heuristic cluster-based confidence metrics and post-selection strategies based on them. (a) Given detector outcomes, a clustering-based decoder constructs valid clusters and performs decoding within each cluster in parallel. As an example, we show a distance-9 surface code patch under bit-flip noise. Blue squares denote violated detectors (check nodes), red circles indicate fault nodes within clusters grown by the decoder, and circles marked with $\times$’ represent the final corrections obtained by decoding each cluster. Note that the depicted cluster configuration is provided solely for illustration; actual results may vary depending on the decoder. (b) Cluster statistics are then used to quantify decoding confidence. As the size and number of clusters grow, a larger portion of the logical operator’s support tends to overlap with clusters, thereby increasing the probability of logical failure. Based on this intuition, we define two metrics (the cluster size and LLR norm fractions) in Definitions 1 and 2, and introduce post-selection strategies that abort a trial when the metric value exceeds a chosen cutoff. (c) This approach is further extended to real-time decoding with the sliding-window framework. After decoding each window of $W$ rounds and committing its first $F$ rounds, clusters are constructed from committed errors within the latest LFLF rounds (where $L$ denotes the lookback window size), and a metric is evaluated based on them. If the value exceeds the cutoff, the trial is immediately aborted, thereby avoiding unnecessary computation.
As illustrated in Fig. 1(a), a clustering-based decoder first identifies valid clusters according to a specific rule, which together can explain all violated detectors. These clusters are then decoded individually to produce local corrections, which are then combined into the global correction.
To define terms more rigorously, the check matrix is a binary matrix indicating how error mechanisms (columns) and detectors (rows) are connected, which forms the biadjacency matrix of the bipartite Tanner graph with check nodes (corresponding to detectors) and fault nodes (corresponding to error mechanisms). The fault graph is constructed by projecting the Tanner graph onto fault nodes; namely, a pair of its vertices is connected within the fault graph if the corresponding error mechanisms are commonly involved in at least one detector. A cluster is a set of error mechanisms that forms a connected region in the fault graph. It is considered valid if it contains at least one solution consistent with all detectors that nontrivially involve any error mechanisms in the cluster.
The union-find (UF) decoder is a representative example of this approach applicable to topological codes [30], which works by growing and fusing clusters incrementally until becoming valid (by checking parity), followed by identifying a spanning tree for each cluster (the peeling decoder). The UF decoder was later generalized to quantum low-density parity-check (QLDPC) codes [31]. Another promising strategy involves forming clusters guided by BP outcomes (such that high-likelihood error mechanisms from BP are preferentially included in clusters), which underlies decoders such as BP plus localized statistics decoding (BP+LSD) [32] and ambiguity clustering (AC) [33].
Specifically, the BP+LSD decoder [32] first runs BP, accepting its solution if it converges, and otherwise applying LSD postprocessing. In LSD, clusters are grown incrementally (prioritizing high-likelihood error mechanisms) until they become valid. Each cluster is then decoded by inverting the submatrix of the check matrix restricted to the relevant detectors and a set of linearly independent columns with highest BP likelihoods within the cluster. Crucially, a matrix factorization algorithm called on-the-fly elimination enables efficient validity check and local decoding; since this can be done in parallel on multiple small clusters, BP+LSD can achieve reduced runtime compared to the BP plus ordered statistics decoding (BP+OSD) decoder [51, 49]
2.2.2 Heuristic confidence metrics from cluster statistics
We construct heuristic confidence metrics by leveraging cluster statistics obtained from clustering-based decoders, such as cluster sizes and aggregated error probabilities within clusters. Intuitively, smaller error clusters indicate better-localized error configurations in terms of the connectivity structure on the Tanner graph, resulting in less ambiguity during error inference. As the size and number of clusters increase, the portion of a logical operator’s support that lies outside the clusters will shrink, and therefore the probability of decoding failure (i.e., the probability that a logical error occurs due to an error chain existing outside the clusters) will increase; see Fig. 1(b).
Based on these intuitions, we propose two families of heuristic confidence metrics (strictly, “inverse” confidence metrics), parametrized by a positive number $\alpha$ (that can be infinity $\infty$): the cluster size $\alpha$-norm fraction and the cluster LLR $\alpha$-norm fraction, defined as follows:
Definition 1 (Cluster size $\alpha$-norm fraction).
For a particular set of error mechanisms $\mathcal{E}$, clusters $\quantity{C_{i}}_{i}$ (such that $\forall i,C_{i}\subseteq\mathcal{E}_{i}$), and a positive real number $\alpha$, the cluster size $\alpha$-norm fraction is defined as
[…]
The $\alpha=\infty$ case is additionally defined as
[…]
Definition 2 (Cluster LLR $\alpha$-norm fraction).
For a particular set of error mechanisms $\mathcal{E}$, clusters $\quantity{C_{i}}_{i}$ (such that $\forall i,C_{i}\subseteq\mathcal{E}_{i}$), and a positive real number $\alpha$, the cluster LLR $\alpha$-norm fraction is defined as
$Q_{\mathrm{LLR}}^{(\alpha)}\quantity(\quantity{C_{i}}_{i};\mathcal{E}):=\frac{1}{\sum_{e\in\mathcal{E}}w_{e}}\quantity[\sum_{i}\quantity(\sum_{e\in C_{i}}w_{e})^{\alpha}]^{1/\alpha}\in[0,1],$
where $w_{e}:=\log[(1-p_{e})/p_{e}]$ is the prior LLR of error mechanism $e$. The $\alpha=\infty$ case is additionally defined as
$Q_{\mathrm{LLR}}^{(\infty)}\quantity(\quantity{C_{i}}_{i};\mathcal{E}):=\frac{1}{\sum_{e\in\mathcal{E}}w_{e}}\max_{i}\sum_{e\in C_{i}}w_{e}\in[0,1].$
These metrics are expected to correlate inversely with decoding confidence. The cluster LLR $\alpha$-norm fraction ($Q_{\mathrm{LLR}}^{(\alpha)}$) additionally weights errors by their LLRs, reflecting that clusters with lower LLR values (higher error probabilities) are more likely to contain real errors. The parameter $\alpha$ controls the influence of large clusters: As $\alpha$ increases, larger clusters dominate the metric, with the limiting case $\alpha=\infty$ considering only the largest cluster.
The set of error mechanisms $\mathcal{E}$ in the definitions may be the full error mechanism set of a DEM, or a subset when disjoint groups of mechanisms are decoded independently. For instance, in Calderbank–Shor–Steane (CSS) codes, if circuits involve only logical $Z$ resets and measurements, it suffices to take $\mathcal{E}$ as the set of $X$ errors to capture confidence information, provided that correlations between $X$ and $Z$ errors are ignored. This approach prevents errors unrelated to an observable from affecting its confidence evaluation.
Note that previous works [29, 17] have proposed cluster-based confidence metrics as well, based on the weight of the shortest segment among all nontrivial error strings that does not overlap with any cluster. Our approach differs by relying solely on information intrinsic to each cluster, without requiring explicit knowledge of the supports of logical operators (whose exhaustive search can be computationally expensive for general codes).
2.2.3 Global post-selection strategy
Based on the cluster norm fraction metrics in Definitions 1 and 2, we propose the following post-selection strategy (the term “global decoding” is used to distinguish from the real-time strategy):
Strategy 1 (Cluster-based post-selection for global decoding).
Execute a clustering-based decoder and collect cluster size or LLR data (optionally restricted to error mechanisms involving the observables to be corrected). Given a method $m\in\quantity{\mathrm{size},\mathrm{LLR}}$, norm order $\alpha>0$, and cutoff threshold $c\in[0,1]$, accept the decoding result only when $Q_{m}^{(\alpha)}\leq c$.
This strategy offers several advantages over the logical gap method:
  1. 1.
    Some clustering-based decoders such as the generalized UF, BP+LSD, and AC decoders are applicable to general QLDPC codes.
  2. 2.
    No comparative decoding is required; a single decoder execution suffices to compute the metrics, while computing the logical gap requires $2^{k}$ executions for $k$ logical observables.
  3. 3.
    The approach integrates naturally with real-time modular decoding (e.g., sliding-window methods [19, 40, 41, 42, 43, 44, 45, 32, 46]), as global cluster statistics can be inferred from partial information in intermediate decoding outcomes.
  4. 4.
    Abort decisions can be made after constructing valid clusters but before running local decoders, avoiding unnecessary computational overheads when aborting. However, this advantage may be modest since cluster construction is often more computationally heavy than its decoding. For instance, the BP+LSD decoder requires $O(n^{3})$ basic operations to build a valid cluster while $O(n^{2})$ to decode it, where $n$ is the cluster size [32].
2.3 Real-time post-selection strategy
While global decoders that process all syndrome data at once are useful for proof-of-principle QEC analysis, they can be impractical for real quantum computing systems. In practice, feedforward corrections must be computed in real time before executing any non-Clifford gate, as such corrections may involve Clifford operations that cannot be absorbed into classical Pauli frame updates. This requirement necessiates real-time decoding, which needs to be sufficiently fast to prevent the backlog problem that could lead to an exponential slowdown during computation [1, 40, 63, 64].
We extend our post-selection strategy to a real-time setting by integrating it with the sliding-window decoding framework for general QLDPC codes [42, 43, 44, 45, 32, 46]. Rather than collecting clusters statistics from global decoding outcomes, we rely only on data from a constant number of recent windows as shown in Fig. 1(c), enabling us to compute a cluster-confidence metric in real time and to triger early aborts when necessary.
The sliding window decoding framework decomposes the overall decoding problem into smaller, overlapping sub-problems defined on “windows”, which can be solved incrementally. By partitioning along the time direction, a partial correction can be derived from the syndrome data collected so far, enabling real-time decoding. The algorithm is characterized by two positive integers $(W,F)$: the window size $W$ (which governs decoding speed and needs to be small enough to avoid the backlog problem) and the commit size $F$ (which corresponds to the width of overlapping regions between adjacent windows). See the Methods section for more details on the algorithm.
To integrate our post-selection strategy with the sliding-window framework, we employ a clustering-based decoder as the inner decoder. As described in Fig. 1(c), we track fractions of clusters overlapping with committed regions and evaluate a cluster norm fraction metric based on the $L$ most recent committed regions. Importantly, whether to abort is judged after decoding each window, potentially reducing the retry cost. The detailed strategy is as follows:
Strategy 2 (Cluster-based post-selection for real-time sliding-window decoding).
Let $L\geq 1$ be the lookback window size, $m\in\{\mathrm{size},\mathrm{LLR}\}$ the cluster weighting method, $\alpha>0$ the norm order, and $c\in[0,1]$ the cutoff threshold. After decoding each window $w\geq L-1$:
  1. 1.
    Record the clusters $\quantity{C_{w,i}}_{i}$ obtained from the current window.
  2. 2.
    Identify all committed error mechanisms that lie within clusters from the last $L$ windows:
    $\mathcal{E}_{L,w}^{\mathrm{inside}}\coloneqq\bigcup_{w^{\prime}=w-L+1}^{w}\quantity(\bigcup_{i}C_{w^{\prime},i}\cap\mathcal{E}_{w^{\prime}}^{\text{commit}}).$
    Here, $\mathcal{E}_{w}^{\text{commit}}$ is defined as the set of error mechanisms committed from window $w$; see Eq. (2) in the Methods section for the formal definition.
  3. 3.
    Construct “committed clusters” $\{C_{L,w,i}^{\mathrm{commit}}\}_{i}$ by grouping the error mechanisms in $\mathcal{E}_{L,w}^{\mathrm{inside}}$ into connected components within the fault graph.
  4. 4.
    Compute the cluster norm fraction
    $Q\coloneqq Q_{m}^{(\alpha)}\quantity(\quantity{C_{L,w,i}^{\mathrm{commit}}}_{i};\bigcup_{w^{\prime}=w-L+1}^{w}\mathcal{E}_{w^{\prime}}^{\text{commit}})$
    for these committed clusters.
  5. 5.
    If $Q>c$, abort the trial immediately.
  6. 6.
    Otherwise, proceed to decode the next window and repeat this process until either it is aborted or the circuit ends.
2.4 Performance analysis
We present a comprehensive numerical evaluation of the post-selection strategies introduced in Strategies 1 and 2. Our simulations employ the BP+LSD-0 decoder [32] for collecting cluster statistics, which is chosen for its broad applicability to arbitrary QLDPC codes, its error-correcting performance comparable to the BP+OSD decoder [51, 49], and its availability through the open-source Python package ldpc [65]. To ensure cluster statistics are available for all samples, we modified the decoder to execute the LSD subroutine even when BP converges successfully. All simulations use 30 BP iterations with the min-sum method; further optimization of these parameters could potentially enhance the numerical results. Details on our simulation methods are presented in the Methods section.
2.4.1 Global strategy analysis
We conduct numerical simulations using memory experiments with $T=d$ rounds of syndrome extraction for logical $Z$ observables. Our analysis covers three representative code families: rotated surface codes, bivariate bicycle (BB) codes [35, 36], and hypergraph product (HGP) codes [37, 38, 39], with code distance $d$.
The circuits are subjected to the uniform circuit-level noise model with error strength $p$, where each single-qubit gate, two-qubit gate, or idling is followed by depolarizing errors with probability $p$, and resets/measurements are flipped with the same probability. After decoding, we compute cluster size and LLR norm fractions with norm orders $\alpha\in\quantity{0.5,1,2,\infty}$, restricted to error mechanisms associated with $Z$-type detectors. Strategy 1 is used for determining acceptance, and the logical error rate $p_{\mathrm{log}}$ (at which any logical observable is erroneous) is evaluated over accepted samples.
Our primary focus is quantifying the trade-off between $p_{\mathrm{log}}$ and the abort rate $p_{\mathrm{abort}}$: How much can the logical error rate be reduced by aborting a certain fraction of low-confidence samples? For comparison, we consider two other easy-to-compute baseline metrics:
  • Correction weight: The log-likelihood weight of the decoder’s correction, as defined in Eq. (1).
  • Detector density: The fraction of violated detectors relative to the total detector count, which generalizes the “non-equilibrium magnetization” heuristic from Ref. [28].
Both baseline metrics have inverse correlation with decoder confidence. Notably, detector density offers exceptional computational efficiency as it can be computed before decoder execution, unlike other post-decoding confidence metrics. For surface codes, we additionally evaluate the logical gap from the MWPM decoder.
Figure 2: Post-selection analysis of global decoding for rotated surface codes. (a) Logical error rates $p_{\mathrm{log}}$ are plotted against $p_{\mathrm{abort}}$ for Strategy 1 based on the cluster LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$, across different physical error rates $p\in\{0.001,0.003,0.005,0.01\}$ and code distances $d\in\{5,9,13\}$. The values of $p_{\mathrm{log}}$ without post-selection ($p_{\mathrm{abort}}=0$) are emphasized as filled circles. For $p=0.001$, an inset displaying the same data with a logarithmic scale on $p_{\mathrm{abort}}$ is included. Shaded regions indicate 95% confidence intervals. (b) Our strategy based on $Q_{\mathrm{LLR}}^{(2)}$ is compared with three other baseline strategies (the logical gap calculated by MWPM, correction weight, and detector density) for $d=13$ and $p=0.005$. (c) Various strategies are compared at a fixed abort rate $p_{\mathrm{abort}}=0.3$. We consider four norm orders $\alpha\in\{0.5,1,2,\infty\}$ (specified above or below the markers) for the cluster size and LLR norm fractions. (d) Required abort rates $p_{\mathrm{abort}}$ to achieve target values of $p_{\mathrm{log}}\in\{10^{-3},10^{-6}\}$ are plotted against $p$ when using the metric $Q_{\mathrm{LLR}}^{(2)}$.
Surface codes.
Figure 2 presents comprehensive results for surface codes with distances $d\in\quantity{5,9,13}$. In Fig. 2(a), we plot the trade-off between $p_{\mathrm{log}}$ and $p_{\mathrm{abort}}$ using the cluster LLR 2-norm fraction ($Q_{\mathrm{LLR}}^{(2)}$) across physical error rates $p\in\quantity{0.001,0.003,0.005,0.01}$. Note that the subplot for $p=0.001$ includes an inset displaying the same data with a logarithmic scale on $p_{\mathrm{abort}}$, as $p_{\mathrm{log}}$ varies rapidly near $p_{\mathrm{abort}}=0$. For most parameter regimes (particularly when $p\leq 0.005$), the strategy shows several orders of magnitude improvement in $p_{\mathrm{log}}$ with modest abort rates.
In Fig. 2(b), we benchmark our cluster-based approach against established methods for $(d,p)=(13,0.005)$, comparing $Q_{\mathrm{LLR}}^{(2)}$ with the logical gap (computed via MWPM), correction weight, and detector density. As expected, our method’s performance falls between the optimal logical gap strategy and the simpler baseline metrics.
Figure 2(c) also compares various strategies, but at a fixed abort rate $p_{\mathrm{abort}}=0.3$, with different values of the norm order $\alpha$ (specified above or below each data point) plotted separately for cluster-based metrics. The results show that the choice of $\alpha$ has little impact on performance as long as $\alpha\geq 1$. Lastly, Fig. 2(d) examines the question of how much we need to abort to achieve a target $p_{\mathrm{log}}$, which may be more practically relevant. The required abort rates to reach target values $p_{\mathrm{log}}\in\quantity{10^{-3},10^{-6}}$ via the strategy based on $Q_{\mathrm{LLR}}^{(2)}$ are plotted against $p$ across different code distances.
Figure 3: Post-selection analysis of global decoding for bivariate bicycle codes. Two variants of bivariate bicycle codes are considered: $[[144,12,12]]$ and $[[72,12,6]]$. Shaded regions represent 95% confidence intervals. (a) Any-observable logical error rates $p_{\mathrm{log}}$ are plotted against the abort rate $p_{\mathrm{abort}}$ at $p\in\{0.001,0.003,0.005\}$ for Strategy 1 based on $Q_{\mathrm{LLR}}^{(2)}$. For $p=0.001$, a log-scale inset is included. (b), (c) Various strategies are compared for the $[[144,12,12]]$ BB code at $p=0.003$. $p_{\mathrm{abort}}$ is fixed to 0.3 in (c). (d) Required $p_{\mathrm{abort}}$ to achieve target values of $p_{\mathrm{log}}\in\{10^{-3},10^{-6}\}$ are plotted against pp.
Bivariate bicycle codes.
Figure 3 presents the simulation results for two instances of BB codes: $[[72,12,6]]$ and $[[144,12,12]]$ [36]. (Here, $[[n,k,d]]$ characterizes a code with $n$ physical qubits, $k$ logial qubits, and the code distance $d$.) In Fig. 3(a), $p_{\mathrm{log}}$ is plotted against $p_{\mathrm{abort}}$ at $p\in\quantity{0.001,0.003,0.005}$, where the subplot for $p=0.001$ includes a log-scale inset. Various strategies are compared in Fig. 3(b) and (c) for the $[[144,12,12]]$ code at $p=0.003$, confirming that the performance of the cluster-based strategy surpasses the correction weight strategy by up to around two orders of magnitude and does not severely depend on the norm order $\alpha$ as long as $\alpha\geq 1$. The required abort rates to reach target values $p_{\mathrm{log}}\in\quantity{10^{-3},10^{-6}}$ are plotted against $p$ in Fig. 3(d).
Figure 4: Post-selection analysis of global decoding for a hypergraph product code. We consider a $[[225,9,6]]$ $(3,4)$-regular hypergraph product code, defined from the product of a 12-variable classical LDPC codes with itself. (a) Any-observable logical error rates $p_{\mathrm{log}}$ are plotted against the abort rate $p_{\mathrm{abort}}$ at $p=0.001$, comparing Strategy 1 based on $Q_{\mathrm{LLR}}^{(2)}$ with the baseline metrics. Shaded regions represent 95% confidence intervals. (b) Various strategies are compared at a fixed $p_{\mathrm{abort}}=0.3$.
Hypergraph product codes.
We evaluate our approach on a $(3,4)$-regular HGP code with parameters $[[225,9,6]]$, constructed by taking the product of a 12-variable classical LDPC code with itself, whose check matrix is given in the Methods section. This particular code offers practical advantages with its depth-8 syndrome extraction circuit [46], which we employ in our simulations. Figure 4 shows the results at $p=0.001$, where our cluster-based approach achieves approximately two orders of magnitude improvement in $p_{\mathrm{log}}$ compared to baseline methods, which is consistent with the performance gains observed for BB codes.
Additional simulation results are provided in the Supplementary Material, including comprehensive data for all parameter combinations (Supplementary Figures 1–3), a detailed examination of the influence of the norm order $\alpha$ (Supplementary Figure 4), and an evaluation of the alternative “conv-max-conf” setting, which treats BP-converged samples as having maximum decoding confidence (Supplementary Figure 5). The conv-max-conf setting avoids the need to modify BP+LSD to enforce LSD execution regardless of BP convergence, but is found to degrade post-selection performance in most cases.
2.4.2 Real-time strategy analysis
We now analyze the real-time post-selection strategy. To fairly assess the impact of “aborting mid-way”, we evaluate the average time cost per accepted shot $\overline{T}_{\mathrm{accepted}}$ instead of $p_{\mathrm{abort}}$, by summing the total number of syndrome extraction rounds elapsed across all samples (including both aborted and accepted runs) and dividing by the number of accepted samples. In other words, $\overline{T}_{\mathrm{accepted}}$ represents the expected time cost required to obtain a single accepted run when immediately retrying after each aborted run. For comparison, global strategies yield $\overline{T}_{\mathrm{accepted}}=T/(1-p_{\mathrm{abort}})$.
Our analysis focuses on $T$-round memory experiments using the surface code with distance $d=13$ and the $[[144,12,12]]$ bivariate bicycle (BB) code. We assume the same circuit-level noise model as the global strategy analysis with $p\in\{0.003,0.005\}$. The decoding is performed using the sliding window framework with window parameters $(W,F)=(5,1)$ for the surface code and $(W,F)=(3,1)$ for the BB code. In both cases, the BP+LSD decoder serves as the inner decoder. For post-selection, we employ the cluster LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$ as our confidence metric with the lookback window size $L\in\{1,2,3,5,7\}$. By varying the cutoff value for this metric, we obtain different trade-offs between $p_{\mathrm{log}}$ and the average time cost per accepted shot $\overline{T}_{\mathrm{accepted}}$.
Figure 5 presents our main results for both codes with $T=d$, clearly exhibiting the trade-off relation between $p_{\mathrm{log}}$ and $\overline{T}_{\mathrm{accepted}}$. For comparison, we include two baseline results: the global post-selection strategy (dashed lines) and standard sliding-window decoding without post-selection ($\times$” marks). Notably, our real-time strategy achieves performance comparable to the global strategy in most cases, and specifically for the BB code at $p=0.005$, it even surpasses global post-selection performance by at most around 1.5 orders of magnitude in $p_{\mathrm{log}}$.
Figure 5: Analysis of real-time post-selection via Strategy 2 for (a) the surface code with $d=13$ and (b) the $[[144,12,12]]$ bivariate bicycle code. Logical error rates $p_{\mathrm{log}}$ are plotted against the average time cost per accepted shots $\overline{T}_{\mathrm{accepted}}$, which represents the average time cost required to succeed when immediately retrying after each aborted attempt. The results are for the memory experiments with $T=d$ at $p\in\{0.003,0.005\}$, decoded with the $(5,1)$ or $(3,1)$ sliding window method. The LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$ is used for the metric and the parameter $L$ of the strategy varies across $\{1,2,3,5,7\}$. For comparison, the values for the global strategy (Strategy 1) and those for sliding window decoding without post-selection are presented additionally as dashed lines and $\times$’ marks, respectively. Shaded regions represent 95% confidence intervals.
Figure 6: Dependence of real-time post-selection performance on the number of rounds $T$ for the $[[144,12,12]]$ bivariate bicycle code. We consider the same setting as Fig. 5 at $p=0.003$ and $L=3$ with varying $T\in\{6,9,\cdots,24\}$. All error bars represent 95% confidence intervals. (a) The logical error rate per round ($p_{\mathrm{log}}/T$) and acceptance rate ($1-p_{\mathrm{abort}}$) are plotted against $T$ for various cutoff values as well as the baseline scenario without post-selection. Three representative cutoff values (specified in the legend) and the “no post-selection” scenario are highlighted with bold lines and bigger markers. Both vertical axes are on a logarithmic scale, and the solid lines represent linear fits on this scale, showing that $p_{\mathrm{log}}/T$ is almost constant while $\log(1-p_{\mathrm{abort}})$ is nearly linear on $T$. (b) $p_{\mathrm{log}}/T$ at the largest $T=24$ and the abort rate per round $p_{\mathrm{abort}}^{\mathrm{round}}$ are plotted against the cutoff value. Here, $p_{\mathrm{abort}}^{\mathrm{round}}\coloneqq 1-1/\exp(G)$, where $G$ is the slope of the linear fit of $\log(1-p_{\mathrm{abort}})$ against $T$ in (a). Solid and dashed lines represent linear fits on a log-log scale. Note that the linear fits are placed just for visual aid and may not be valid beyond the presented region. Specifically, the “no post-selection” scenario ($c=1$) has $p_{\mathrm{log}}/T\approx 2.6\times 10^{-3}$ (blue dotted horizontal line) and $p_{\mathrm{abort}}^{\mathrm{round}}=0$, which deviate significantly from the lines. (c) Dependence of the global strategy performance on $T$ with the cutoff fixed to $0.004$ is presented for comparison. Unlike the real-time strategy, $p_{\mathrm{log}}/T$ varies on $T$.
The real-time strategy can be particularly useful for circuits involving many syndrome extraction rounds, where the global strategy may become computationally prohibitive. To assess the practical viability of our approach in such scenarios, we investigate how its performance scales with the total number of rounds $T$. Figure 6(a) examines the scaling behavior by plotting both $p_{\mathrm{log}}/T$ and the acceptance rate $1-p_{\mathrm{abort}}$ (in a logarithmic scale) against $T$ for the $[[144,12,12]]$ BB code under various fixed cutoff values, including the scenario without post-selection as a baseline. (Note that we here plot the acceptance rate rather than $\overline{T}_{\mathrm{accepted}}$, as it directly reveals its decay over rounds and allows us to estimate the per-round abort rate.)
Two key observations emerge from Fig. 6(a): First, $p_{\mathrm{log}}/T$ remains nearly constant across different $T$ values even under post-selection, which indicates that logical errors accumulate linearly with the number of rounds, preserving the expected scaling behavior shown in the case without post-selection. Second, $\log(1-p_{\mathrm{abort}})$ exhibits clear linear dependence on $T$. This behavior can be interpreted as every single round has an equal per-round abort rate $p_{\mathrm{abort}}^{\mathrm{round}}$, which is related to $p_{\mathrm{abort}}$ as $p_{\mathrm{abort}}=1-(1-p_{\mathrm{abort}}^{\mathrm{round}})^{T}$ and thus can be estimated from the slope of the fit in the right panel. In Fig. 6(b), we plot the estimated values of $p_{\mathrm{log}}/T$ (at $T=24$) and $p_{\mathrm{abort}}^{\mathrm{round}}$ against the cutoff, clearly illustrating the trade-off relation between them.
To highlight the advantages of the real-time approach, we conduct a parallel analysis for the global strategy in Fig. 6(c) with a fixed cutoff of $c=0.004$. Crucially, $p_{\mathrm{log}}/T$ is no longer invariant under $T$, underscoring the practical necessity of the real-time strategy particularly for large $T$.
3 Discussion
In this work, we have shown that decoder soft outputs derived from error-cluster structure provide an effective foundation for post-selection in quantum low-density parity check (QLDPC) codes. Leveraging the insight that clustering-based decoders (such as BP+LSD [32]) naturally capture the geometric properties of error configurations in the Tanner graph, we introduced two heuristic confidence metrics based on cluster statistics: cluster size and log-likelihood ratio (LLR) norm fractions. Utilizing these metrics, we developed post-selection strategies that are applicable to general QLDPC codes and require only a single decoding run. This approach offers significant advantages over the conventional logical gap method [15, 16, 17], which is effectively applicable only to a limited family of QEC codes (those compatible with the MWPM decoder, such as surface codes) and suffers from computational overheads that scale exponentially with the number of logical qubits.
Our numerical simulations demonstrated that our post-selection strategies achieve signifcant reductions in the logical error rate $p_{\mathrm{log}}$ (for accepted runs) by several orders of magnitude at moderate abort rates $p_{\mathrm{abort}}$. This performance gain is observed consistently across diverse code families: surface codes, bivariate bicycle codes, and hypergraph product codes (Figs. 24). For instance, by applying our strategy to the $[[144,12,12]]$ bivariate bicycle (BB) code, the logical error rate can be reduced by about three orders of magnitude with the abort rate of only around 1% (19%) at the physical error rate ($p$) of 0.1% (0.3%) under the uniform circuit-level noise model.
Our analysis highlights that post-selection can offer a far more cost-efficient alternative to simply increasing the code distance, particularly when non-determinism is acceptable and the system operates in a slightly sub-threshold regime. For example, at $p=0.1\%$ ($p=0.3\%$), aborting only about 8% (4%) of samples for the $[[72,12,6]]$ BB code achieves a logical error rate comparable to that of the larger $[[144,12,12]]$ BB code without post-selection. In other words, a modest spacetime cost increase of $\times 1.09$ ($\times 1.04$) with post-selection yields an effect roughly equivalent to a $\times 4$ increase in spacetime cost without post-selection.
Furthermore, real-time decoding is a key requirement for practical quantum computing involving non-Clifford gates, which should be sufficiently fast to avoid the backlog problem [1]. We extended our approach to such real-time scenarios through sliding-window decoding, featuring mid-circuit abort decisions that allow further computational cost reduction. The real-time strategy achieves performance comparable to (or even superior to) global post-selection in terms of the average time cost per accepted shot $\overline{T}_{\mathrm{accepted}}$ for achieving a target logical error rate $p_{\mathrm{log}}$ (Fig. 5). Notably, the real-time strategy maintains favorable scaling with the number of rounds $T$, exhibiting nearly constant per-round logical error and abort rates (Fig. 6). In contrast, the global strategy does not preserve this consistency, making the real-time approach especially advantageous in practical regimes with large $T$.
We interpret the gains from cluster-based metrics as evidence that the geometry and concentration of errors (in terms of the connectivity structure on the Tanner graph) are key determinants of decoding ambiguity. Large clusters are inherently problematic because they can span multiple local solutions and logical classes, thereby increasing the likelihood of miscorrection. In contrast, fragmented, smaller clusters present fewer ambiguities and are more readily resolved by the decoder. This geometric perspective explains why cluster norm fractions significantly outperform baseline metrics such as correction weight and detector density, which rely on simple aggregations of error weights or detector flips without capturing the detailed geometric structure of error configurations. Additionally, the performance does not severely depend on the norm order $\alpha$ as long as $\alpha\geq 1$, which implies that, once large clusters dominate the metric, finer details of how residual clusters are formed matter little.
Finally, we identify key limitations and promising directions for future work.
  • Theoretical foundations: Establishing theoretical justifications (e.g., bounds that relate cluster-based metrics to the logical gap) would provide rigorous foundations for these heuristics. Understanding which code or circuit properties (such as expansion [66]) influence post-selection performance would also be valuable.
  • Further performance improvements: While cluster-based metrics offer practical advantages, the logical gap method still outperforms them for surface codes. Developing better metrics that approach logical gap performance while retaining computational efficiency and generalizability remains an open challenge.
  • Broader noise models: Testing our strategies beyond standard circuit-level noise (including strongly biased, correlated, or other realistic noise models) would demonstrate their robustness. In these cases, cluster size and LLR norm fractions may lead to more pronounced differences in post-selection performance compared to our current settings.
  • Practical applications: Evaluating the effectiveness of these strategies for specific quantum computing tasks, such as magic state preparation, would verify their practical utility.
4 Methods
4.1 Sliding-window decoding
We here describe the sliding-window decoding framework that we use for our real-time post-selection. The framework with the window size $W$ and the commit size $F$ works as follows: The set of detectors is partitioned into a sequence of temporal windows, each spanning $W$ rounds, such that two consecutive windows overlap by WFW-F rounds. Within each window, a decoder identifies a correction; however, only a fraction of the correction associated with the first $F$ rounds of each window are committed to the global solution. The syndrome is updated accordingly, and the decoder proceeds to the next window. This process repeats until all detector rounds have been processed.
We now describe the algorithm more formally. Consider a check matrix $H\in\{0,1\}^{r\times N}$ for $r$ detectors and $N$ error mechanisms. Let $\mathbf{s}\in\{0,1\}^{r}$ denote the syndrome vector and $\mathbf{p}\in[0,1]^{N}$ the prior probability vector for error mechanisms. We denote the sets of detector and error mechanism indices as $\mathcal{D}\coloneqq\{1,\cdots,r\}$ and $\mathcal{E}\coloneqq\{1,\cdots,N\}$, respectively. Each detector $i\in\mathcal{D}$ is assigned a time coordinate $t_{i}\in\{0,1,2,\cdots\}$. The algorithm maintains a global correction vector $\hat{\mathbf{e}}$, initially set to $\hat{\mathbf{e}}\leftarrow\mathbf{0}=(0,\cdots,0)\in\{0,1\}^{N}$.
For each window $w\in\{0,1,2,\cdots\}$, the corresponding detector subset is defined as
$\mathcal{D}_{w}\coloneqq\quantity{i\in\mathcal{D}:wF\leq t_{i}\leq wF+W-1}$.
To construct the decoding sub-problem for window $w$, we identify relevant error mechanisms and detectors. First, we determine the “active error mechanisms” for this window, which are error mechanisms that connect to detectors within the current window, excluding any that have already been committed in previous windows. Formally, denoting the set of error mechanisms commited so far as $\mathcal{C}$ (initially empty), the set of active error mechanisms is given as
$\mathcal{E}_{w}\coloneqq\quantity{j\in\mathcal{E}:\exists i\in\mathcal{D}_{w}\text{ s.t. }H[i,j]=1}\setminus\mathcal{C},$
where $H[i,j]$ denotes the $(i,j)$ element of $H$.
The decoding sub-problem for window $w$ is then defined by the sub-matrix $H_{w}=H[\mathcal{D}_{w},\mathcal{E}_{w}]$, the reduced syndrome vector $\mathbf{s}_{w}=\mathbf{s}[\mathcal{D}_{w}]$, and the reduced prior probability vector $\mathbf{p}_{w}=\mathbf{p}[\mathcal{E}_{w}]$. This sub-problem can be solved using any chosen inner decoder.
After obtaining the window solution $\hat{\mathbf{e}}_{w}$ (satisfying $H_{w}\hat{\mathbf{e}}_{w}=\mathbf{s}_{w}$), only error mechanisms involved in detectors within the first $F$ rounds are committed:
$\mathcal{E}_{w}^{\text{commit}}\coloneqq\quantity{j\in\mathcal{E}_{w}:\exists i\in\mathcal{D}_{w}\text{ s.t. }wF\leq t_{i}\leq wF+F-1\text{ and }H_{ij}=1}$ (2)
Next, the global state is updated by incorporating the committed corrections. To map the window-local solution back to the global problem, we define a commit vector $\hat{\mathbf{e}}_{w}^{\mathrm{commit}}\in\{0,1\}^{N}$ as
$\hat{\mathbf{e}}_{w}^{\mathrm{commit}}[j]=\begin{cases}\hat{\mathbf{e}}_{w}[\phi_{w}(j)]&\text{if }j\in\mathcal{E}_{w}^{\text{commit}},\\ 0&\text{otherwise},\end{cases}$
for each $j\in\mathcal{E}$, where $\phi_{w}$ maps global error mechanism indices to the corresponding local indices in $\mathcal{E}_{w}$.
The global correction vector and syndrome are then updated as $\hat{\mathbf{e}}\leftarrow\hat{\mathbf{e}}\oplus\hat{\mathbf{e}}_{w}^{\mathrm{commit}}$ and $\mathbf{s}\leftarrow\mathbf{s}\oplus H\hat{\mathbf{e}}_{w}^{\mathrm{commit}}$, where $\oplus$” denotes addition modulo 2. This update effectively removes the syndrome contributions from committed error corrections. Finally, the set of committed error mechanisms is expanded: $\mathcal{C}\leftarrow\mathcal{C}\cup\mathcal{E}_{w}^{\text{commit}}$.
The algorithm continues until the windows decoded so far cover all detectors (i.e., $wF+W-1\geq\max\{t_{i}:i\in\mathcal{D}\}$). In the final window $w_{\mathrm{final}}$, all remaining active error mechanisms are committed: $\mathcal{E}_{w_{\mathrm{final}}}^{\text{commit}}\coloneqq\mathcal{E}_{w_{\mathrm{final}}}$ instead of Eq. (2).
4.2 Technical details on the simulation methods
We here provide technical details on our numerical simulation methods used to obtain the data in the Results section.
4.2.1 Circuit generation and sampling
All simulations utilize the stim library [47] to sample circuit measurement outcomes. We generate the memory circuits for different QEC codes as follows:
  • Surface codes: Generated using stim.Circuit.generated("surface_code:rotated_memory_z").
  • BB codes: Generated using the code provided in Ref. [43].
  • (3, 4)-regular HGP code: Generated using the QUITS library [46] with the setting HgpCode(h,h).build_graph(seed=22), where h represents the following check matrix of the classical code that forms the HGP code:
[…]
    This specific $[[225,9,6]]$ HGP code is identical to the one optimized in Ref. [46], featuring a depth-8 syndrome extraction circuit.
4.2.2 Noise model
We apply a uniform circuit-level depolarizing noise model characterized by an error rate parameter $p$:
  • Single-qubit Clifford gates are followed by depolarizing errors with probability $p$ (i.e., each of $X$, $Y$, and $Z$ occurs with probability $p/3$).
  • Two-qubit Clifford gates are followed by two-qubit depolarizing errors with probability $p$ (i.e., each of the 15 nontrivial two-qubit Pauli errors occurs with probability $p/15$).
  • Idling qubits between consecutive time steps are subject to depolarizing errors with probability $p$.
  • Resets in the $Z$ ($X$) basis are followed by an $X$ ($Z$) error with probability $p$.
  • Measurement outcomes are randomly flipped with probability $p$.
4.2.3 Decoder implementation
While our post-selection strategies can be applied to any clustering-based decoder, we specifically employ the BP+LSD decoder [32] for our simulations, implemented in the ldpc library [65].
The original decoder implementation skips LSD postprocessing when BP converges, which makes it impossible to obtain cluster-based metrics in converged cases. To address this limitation, we modified the decoder to execute LSD regardless of BP convergence, and this modification has been integrated into the official ldpc library. Additionally, we test our strategies under the alternative assumption that BP convergence indicates maximum confidence; these results are presented in the supplementary information.
We use the following parameters for BP+LSD across all simulations:
  • max_iter: 30 (Same setting as in Ref. [32])
  • bp_method: "minimum_sum"
  • lsd_method: "LSD_0"
4.2.4 Overall pipeline implementation
The complete simulation pipeline is implemented in our GitHub repository ldpc-post-selection [^1] https://github.com/seokhyung-lee/ldpc-post-selection. The repository includes a SoftOutputsBpLsdDecoder class that supports executing the modified BP+LSD decoder (with optional sliding-window integration) and extracting both corrections and soft outputs. Detailed usage instructions are provided in the repository documentation.
Data availability
The aggregated data that support the findings of this study are openly available in Harvard Dataverse at DOI:10.7910/DVN/MNYLTA. The raw data are not uploaded due to their size but are available upon reasonable request to the authors.
Code availability
The source codes used for the simulations in this work, including the ldpc-post-selection package, are available in GitHub: https://github.com/seokhyung-lee/ldpc-post-selection.
Acknowledgements
We thank Timo Hillmann, Nicholas Fazio, Dominic Williamson, Mingyu Kang, and Hyukgun Kwon for helpful discussions and comments. We also thank Joschka Roffe for integrating our BP+LSD customization into the official ldpc library.
Funding
This work is supported by the Australian Research Council via the Centre of Excellence in Engineered Quantum Systems (EQUS) Project No. CE170100009, and by the Intelligence Advanced Research Projects Activity (IARPA) through the Entangled Logical Qubits program Cooperative Agreement Number W911NF-23-2-0223.
Author contributions
S.H.L. conceived the idea and all authors contributed to developing it. S.H.L. conducted simulations and all authors interpreted the results together. S.H.L. mainly wrote the manuscript and L.E. made the first draft of Figure 1. All authors reviewed the manuscript.
Competing interests
The authors declare no competing interests.